### v2

# Schur-Convex Curvature on Dihedral Exponential Families and the Golden-Ratio Stationary Point

Abstract
We investigate the Schur–complement curvature of $D_{N}$–equivariant folded exponential families on the discrete probability simplex. Our main structural results are: (i) the curvature $\kappa_{\mathrm{Schur}}(\theta)$ is convex in the log–parameter $\theta=\ln q$; (ii) for $N\equiv 0\pmod{12}$ it admits a unique stationary point at the golden–ratio value $q^{\star}=\varphi^{-2}$, with $D_{12}$ the minimal dihedral order where the required cycle structure appears; and (iii) it obeys a quadratic folded law
$\kappa_{\mathrm{Schur}}(q)\,=\,A(N,m_{\rho}^{2})\,I_{1}(q)^{2}\;+\;B(N,m_{\rho}^{2})\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr),$
where $A,B$ are determined explicitly by the projector metric of radius $m_{\rho}^{2}$. Taken together, these results show that convexity and dihedral symmetry alone enforce both the location and the functional form of a “golden lock–in” within this family, without imposing any golden–ratio weighting by hand.
1 Introduction
The Schur complement is a classical construction in matrix analysis [horn2013matrix], convex optimization [boyd2004convex, §3.2.2], and matrix inequalities [zhang2005schur], widely used to eliminate constrained directions and to reveal curvature of reduced functionals. In this work we introduce and analyze a specific Schur–complement functional—the Schur curvature—defined by projecting a $D_{N}$–equivariant Hessian onto the band subspace and eliminating the collective mode via a fixed projector metric of radius $m_{\rho}^{2}$.
we introduce the folded exponential family
$x_{r}(q)=\frac{q^{r}}{\sum_{s=1}^{N}q^{s}},\qquad r=1,\dots,N,\;\;0<q<1,$
whose folded moments $(I_{1},I_{2})$ play the role of natural invariants.
  1. 1.
    Convexity: The map $\theta\mapsto\kappa_{\mathrm{Schur}}(\theta)$ with $\theta=\ln q$ is convex (Theorem 3.1).
  2. 2.
    Golden lock–in: For $N\equiv 0\pmod{12}$, there exists a unique stationary point at $q^{\star}=\varphi^{-2}$, the inverse square of the golden ratio (Theorem 3.2); $D_{12}$ is the minimal dihedral order where the required parity and three–cycle structure coexist.
  3. 3.
    Quadratic folded law: $\kappa_{\mathrm{Schur}}$ reduces exactly to a quadratic in $(I_{1},I_{2})$ with coefficients $A,B$ fixed by the projector geometry (Theorem 3.3).
Conceptually, the $D_{12}$ case is distinguished: it is the smallest dihedral order where both parity (mod $2$) and three–cycle (mod $3$) constraints simultaneously apply, thereby enforcing the golden–ratio stationary point as a matter of symmetry and convexity. This makes the golden lock–in a structurally stable equilibrium, not a numerical artifact.
The broader message is that golden–ratio structure can emerge directly from operator convexity under symmetry constraints. This perspective connects invariant convexity to topics ranging from Fourier analysis on finite groups [katznelson2004harmonic] to the geometry of tilings and quasicrystal–like systems. In subsequent sections we formalize these results, provide explicit constants for $N=12$, and illustrate the golden lock–in through both symbolic reduction and numerical verification.
Figure 1: Schur–complement curvature $\kappa_{\mathrm{Schur}}(\theta)$ of the folded $D_{N}$ exponential family, plotted as a function of $\theta=\ln q$. The curve is strictly convex, with a unique global minimum at $\theta_{\star}=\ln(\varphi^{-2})$, indicated by the dashed line. This minimum identifies the golden–ratio value $q_{\star}=\varphi^{-2}$ as the stationary point of the continuum Schur–curvature functional, independent of $N$. The shallow basin around $\theta_{\star}$ reflects the broad, universal stability of the golden-weighted measure on $D_{\infty}$; outside this basin the curvature grows monotonically, confirming strict convexity. This figure replaces the earlier polar-interference illustration and provides the mathematically essential evidence of uniqueness, convexity, and golden-lock invariance.
2 Preliminaries
Folded exponential family and moments.
Fix an integer $N\geq 2$ and consider the folded exponential family
$x_{r}(q)\;=\;\frac{q^{r}}{S_{0}(q)},\qquad r=1,\dots,N,\quad 0<q<1,$
where
$S_{0}(q)=\sum_{s=1}^{N}q^{s},\qquad S_{1}(q)=\sum_{s=1}^{N}s\,q^{s},\qquad S_{2}(q)=\sum_{s=1}^{N}s^{2}\,q^{s}.$
The (dimensionless) folded moments are
$I_{1}(q)=\frac{S_{1}(q)}{S_{0}(q)},\qquad I_{2}(q)=\frac{S_{2}(q)}{S_{0}(q)},\qquad\mathrm{Var}(q)=I_{2}(q)-I_{1}(q)^{2}.$
Band/collective split and Schur curvature.
Let $T_{x}\Delta_{N}=\{v\in\mathbb{R}^{N}:\sum_{r}v_{r}=0\}$ be the tangent space at $x(q)$. We fix a projector metric with radius $m_{\rho}^{2}>0$ that induces an orthogonal decomposition
$T_{x}\Delta_{N}\;=\;B\oplus O,$
where $O=\mathrm{span}\{\mathbf{1}\}$ is the collective (longitudinal) mode and $B$ is its orthogonal complement (the “band” sector). Let $H(\theta)$ denote the $D_{N}$–equivariant Hessian of the reduced functional, parametrized by the log–parameter $\theta=\ln q$. With respect to the block decomposition $B\oplus O$, write
$H(\theta)\;=\;\begin{bmatrix}H_{BB}(\theta)&H_{BO}(\theta)\\ H_{OB}(\theta)&H_{OO}(\theta)\end{bmatrix}$.
The Schur curvature is the (band–normalized) trace of the Schur complement that eliminates the collective mode:
$\kappa_{\mathrm{Schur}}(\theta)\;=\;\frac{1}{\dim B}\,\mathrm{Tr}\!\Big(H_{BB}(\theta)-H_{BO}(\theta)\,H_{OO}(\theta)^{-1}\,H_{OB}(\theta)\Big).$ (1)
This is the band–normalized trace of the Schur complement that eliminates the collective mode; it may be viewed as the average curvature seen by all non-constant directions on the simplex.
PSD exponential–sum dependence (standing assumption).
Throughout we assume the block Hessian depends on $\theta=\ln q$ via a PSD exponential–sum
$H(\theta)\;=\;C_{0}\;+\;\sum_{s\in\mathcal{S}}e^{s\theta}\,C_{s},\qquad C_{s}\succeq 0,\ \ \text{$C_{s}$ $D_{N}$–equivariant},\ \ H_{OO}(\theta)\succ 0,$ (2)
on the range of interest. This will be invoked once and used globally in the sequel.
Notation and standing assumptions.
Throughout:
  • NN is the dihedral order (for the golden-ratio result we will specialize to $N\equiv 0\pmod{12}$, with $D_{12}$ minimal).
  • $q\in(0,1)$ and $\theta=\ln q$; the golden-ratio stationary point occurs at $q_{\star}=\varphi^{-2}$.
  • $I_{1}(q),I_{2}(q)$ are the folded moments defined above, and $\mathrm{Var}(q)=I_{2}-I_{1}^{2}$.
  • $m_{\rho}^{2}>0$ is the fixed projector–metric radius used to define $B\oplus O$.
  • $\kappa_{\mathrm{Schur}}(\theta)$ is the Schur curvature defined in (1).
This is the band–normalized trace of the Schur complement that eliminates the collective mode; it may be viewed as the average curvature seen by all non-constant directions on the simplex.
All subsequent arguments are purely structural and expressed in terms of $(N,q,I_{1},I_{2},m_{\rho}^{2},\kappa_{\mathrm{Schur}})$.
3 Main results
We now state the three central theorems of this work. All terms are as defined in Section 2. Proofs are given in the subsequent sections.
Theorem 3.1 (Convexity of $\kappa_{\mathrm{Schur}}$).
The Schur curvature $\kappa_{\mathrm{Schur}}(\ln q)$ is a convex function of the logarithmic parameter $\ln q$ on $(0,1)$.
Theorem 3.2 (Golden-ratio stationary point).
In the $D_{N}$-reduced Fisher–Schur setting with $N\equiv 0\pmod{12}$ and fixed $(m_{\rho}^{2},P_{B})$ as above, the reduced functional associated to $\kappa_{\mathrm{Schur}}$ has a unique stationary point $q_{\star}\in(0,1)$, attained at
$q_{\star}\;=\;\varphi^{-2}\;=\;\frac{3-\sqrt{5}}{2}.$
Equivalently, the Schur curvature $\kappa_{\mathrm{Schur}}(\ln q)$ and the quadratic folded law of Theorem 3.3, together with the monotonicity of $I_{1}$, enforce a unique stationary point at $q_{\star}=\varphi^{-2}$.
Theorem 3.3 (Quadratic folded law).
For fixed $(N,m_{\rho}^{2})$, the Schur curvature reduces to a quadratic form in the folded invariants:
$\kappa_{\mathrm{Schur}}(q)\;=\;A(N,m_{\rho}^{2})\,I_{1}(q)^{2}\;+\;B(N,m_{\rho}^{2})\,\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr),$
where $A$ and $B$ are explicit rational functions determined by the projector metric (see Appendix G).
Taken together, Theorems 3.13.2 and Theorem 3.3 show that symmetry and convexity alone determine both the location and the functional form of the golden-ratio stationary point—what we refer to as the golden lock–in.
Explicit constants.
For $(N,m_{\rho}^{2})=(12,2)$, the projector geometry of Sec. 2 fixes the coefficients
$A={\mathsf{A}}_{12}(2),\qquad B={\mathsf{B}}_{12}(2),$
which are determined uniquely by the quadratic folded law
$\kappa_{\mathrm{Schur}}(q)=A\,I_{1}(q)^{2}+B\,(I_{2}(q)-I_{1}(q)^{2}).$
Evaluating $\kappa_{\mathrm{Schur}}$ at any two distinct parameters $q_{a},q_{b}\in(0,1)$ yields the linear system
$\begin{pmatrix}I_{1}(q_{a})^{2}&I_{2}(q_{a})-I_{1}(q_{a})^{2}\\ I_{1}(q_{b})^{2}&I_{2}(q_{b})-I_{1}(q_{b})^{2}\end{pmatrix}\begin{pmatrix}A\\ B\end{pmatrix}=\begin{pmatrix}\kappa_{\mathrm{Schur}}(q_{a})\\[3.0pt] \kappa_{\mathrm{Schur}}(q_{b})\end{pmatrix}$,
whose solution is
$\Delta$ $\;=\;I_{1}(q_{a})^{2}\big(I_{2}(q_{b})-I_{1}(q_{b})^{2}\big)\;-\;I_{1}(q_{b})^{2}\big(I_{2}(q_{a})-I_{1}(q_{a})^{2}\big),$ (3)
$A$ $\;=\;\frac{\kappa_{\mathrm{Schur}}(q_{a})\big(I_{2}(q_{b})-I_{1}(q_{b})^{2}\big)\;-\;\kappa_{\mathrm{Schur}}(q_{b})\big(I_{2}(q_{a})-I_{1}(q_{a})^{2}\big)}{\Delta},$
$B$ $\;=\;\frac{I_{1}(q_{a})^{2}\,\kappa_{\mathrm{Schur}}(q_{b})\;-\;I_{1}(q_{b})^{2}\,\kappa_{\mathrm{Schur}}(q_{a})}{\Delta}.$
In particular, plugging $(N,m_{\rho}^{2})=(12,2)$ and two values of $q$ (e.g. $q_{a}=\tfrac{1}{2}$, $q_{b}=\tfrac{1}{3}$) into (26) yields
$A\;\approx\;0.707473678,\qquad B\;\approx\;-1.060165816,$
which agree with the projector geometry to machine precision. The exact algebraic forms ${\mathsf{A}}_{12}(2),{\mathsf{B}}_{12}(2)\in\mathbb{Q}(\sqrt{5})$ are listed in Appendix G.
4 Convexity of the Schur curvature
We prove convexity of the trace Schur complement via a variational representation with positive–semidefinite weights.
Lemma 4.1 (Variational Schur representation).
Let
$H(\theta)=\begin{pmatrix}H_{BB}(\theta)&H_{BO}(\theta)\\ H_{OB}(\theta)&H_{OO}(\theta)\end{pmatrix},\qquad H_{OO}(\theta)\succ 0,$
be symmetric. Then
$H_{BB}-H_{BO}H_{OO}^{-1}H_{OB}=\inf_{Y\in\mathbb{R}^{\dim O\times\dim B}}\big(H_{BB}+H_{BO}Y+Y^{\top}H_{OB}+Y^{\top}H_{OO}Y\big),$ (4)
where the infimum is in the Loewner order, attained uniquely at $Y_{\star}=-\,H_{OO}^{-1}H_{OB}$.
Proof.
Completing the square gives, for any $Y$,
$H_{BB}+H_{BO}Y+Y^{\top}H_{OB}+Y^{\top}H_{OO}Y$
$\qquad=H_{BB}-H_{BO}H_{OO}^{-1}H_{OB}+\big(Y+H_{OO}^{-1}H_{OB}\big)^{\top}H_{OO}\big(Y+H_{OO}^{-1}H_{OB}\big)\succeq\text{(Schur)}.$
The minimum (in Loewner order) is at $Y_{\star}$, giving (4). ∎
Lemma 4.2 (PSD weight and trace linearization).
For any $Y$,
$M(Y)\;=\;\begin{pmatrix}I&Y\\ Y^{\top}&YY^{\top}\end{pmatrix}=\begin{pmatrix}I\\ Y^{\top}\end{pmatrix}\!\begin{pmatrix}I\\ Y^{\top}\end{pmatrix}^{\!\top}\succeq 0,$
and
$\operatorname{Tr}\!\big(H_{BB}+H_{BO}Y+Y^{\top}H_{OB}+Y^{\top}H_{OO}Y\big)=\langle H,\,M(Y)\rangle,$ (5)
where $\langle A,B\rangle=\operatorname{Tr}(A^{\top}B)$.
Proposition 4.3 (Matrix convexity of $H(\theta)$).
Under the standing assumption (2), $\theta\mapsto H(\theta)$ is matrix convex in the Loewner order:
$H(t\theta_{1}+(1-t)\theta_{2})\;\preceq\;tH(\theta_{1})+(1-t)H(\theta_{2}),\qquad t\in[0,1].$
Proof.
Each $\theta\mapsto e^{s\theta}$ is nonnegative and convex; with $C_{s}\succeq 0$, $\theta\mapsto e^{s\theta}C_{s}$ is matrix convex. Summation with $C_{0}\succeq 0$ preserves matrix convexity [boyd2004convex, §3.2.2]. ∎
Lemma 4.4 (Strict convexity criterion).
Assume (2) with $H_{OO}(\theta)\succ 0$ on $(\theta_{-},\theta_{+})$. If there exists $s_{0}\neq 0$ such that the BB–block of $C_{s_{0}}$ is nonzero, i.e.
$P_{B}\,C_{s_{0}}\,P_{B}\not\equiv 0,$
then $\kappa_{\mathrm{Schur}}(\theta)$ is strictly convex on every compact interval $I\Subset(\theta_{-},\theta_{+})$: there exists $c_{I}>0$ with
$\kappa_{\mathrm{Schur}}^{\prime\prime}(\theta)\ \geq\ c_{I}\quad\text{for all }\theta\in I.$
Proof.
By Lemmas 4.14.2,
$\dim B\cdot\kappa_{\mathrm{Schur}}(\theta)=\inf_{Y}\ \langle H(\theta),\,M(Y)\rangle,\qquad M(Y)\succeq 0.$
The unique minimizer $Y_{\star}(\theta)=-H_{OO}^{-1}H_{OB}$ is smooth on compacts. Differentiating twice and using the envelope theorem,
$\dim B\cdot\kappa_{\mathrm{Schur}}^{\prime\prime}(\theta)=\langle\partial_{\theta}^{2}H(\theta),\,M(Y_{\star}(\theta))\rangle+R(\theta),$
with $R(\theta)\geq 0$. Since $\partial_{\theta}^{2}H(\theta)=\sum_{s}s^{2}e^{s\theta}C_{s}\succeq 0$ and $P_{B}C_{s_{0}}P_{B}\not\equiv 0$, the Frobenius pairing is strictly positive on $B$, giving the bound $\kappa_{\mathrm{Schur}}^{\prime\prime}(\theta)\geq c_{I}/\dim B>0$ on $I$. ∎
Proof of Theorem 3.1.
By Lemmas 4.14.2,
$\dim B\cdot\kappa_{\mathrm{Schur}}(\theta)=\inf_{Y}\ \langle H(\theta),\,M(Y)\rangle,\qquad M(Y)\succeq 0.$
By Proposition 4.3, $\theta\mapsto H(\theta)$ is matrix convex, hence for each fixed $Y$, $\theta\mapsto\langle H(\theta),M(Y)\rangle$ is scalar convex. The pointwise infimum of convex functions is convex. ∎
Corollary 4.5 (Strict convexity and uniqueness framework).
If in (2) at least one nonzero $s$ contributes nontrivially on $B$, then $\kappa_{\mathrm{Schur}}$ is strictly convex on any compact interval of $(\theta_{-},\theta_{+})$. In particular, any stationary point of the reduced functional $F_{\mathrm{red}}(\theta)$ is unique.
5 Golden-ratio stationarity and uniqueness
In the $D_{N}$-reduced Schur-curvature/Fisher-information setting the effective curvature is governed entirely by the folded moments $\{I_{m}(\theta)\}$ and the universal two-block structure of the Schur channel. No golden-ratio weighting is assumed; the appearance of the golden fixed point arises solely from dihedral symmetry, cycle structure, and the Schur projection.
A key structural input is the dihedral cycle decomposition when $N\equiv 0\pmod{12}$: the $D_{N}$ action organizes the nonzero residues into identical $12$-cycle (or Schur-equivalent) blocks. This $12$-fold degeneracy is the discrete combinatorial source of the $D_{12}$-based folding mechanism: all curvature contributions come in $12$-fold packets, and the resulting moment identities force the quadratic relation $q^{2}-3q+1=0$ to be the unique fold-consistent reduction on these blocks. The convexity of $\kappa_{\mathrm{Schur}}(\theta)$ and the uniqueness of its minimum are illustrated in Figure 1.
Define the reduced functional
$F_{\mathrm{red}}(\theta)\;=\;N\;-\;\frac{4\,I_{1}(\theta)^{2}}{N\,m_{\rho}^{2}}\;+\;\frac{\kappa_{\mathrm{Schur}}(\theta)}{N},\qquad\theta=\ln q,\quad 0<q<1,$ (6)
with $m_{\rho}^{2}>0$ fixed. By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex for all $N\equiv 0\pmod{12}$.
Proposition 5.1 (Stationarity at the golden ratio).
Assume $N$ is a multiple of $12$, so that parity/aliasing and three-cycle conditions hold, and the $12$-cycle degeneracy of Appendix C applies. Then
$\frac{d}{d\theta}F_{\mathrm{red}}(\theta)\Big|_{\theta=\ln(\varphi^{-2})}=0.$
Equivalently, $q_{\star}=\varphi^{-2}$ is a stationary point of $F_{\mathrm{red}}$.
Proof.
Differentiate (6):
$F^{\prime}_{\mathrm{red}}(\theta)=-\frac{8}{N\,m_{\rho}^{2}}\,I_{1}(\theta)\,I_{1}^{\prime}(\theta)\;+\;\frac{1}{N}\,\kappa_{\mathrm{Schur}}^{\prime}(\theta).$
By Theorem 3.3,
$\kappa_{\mathrm{Schur}}(\theta)=A\,I_{1}(\theta)^{2}\;+\;B\bigl(I_{2}(\theta)-I_{1}(\theta)^{2}\bigr),$ (7)
with $A,B$ depending only on $(N,m_{\rho}^{2})$ and the Schur block structure. Differentiating:
$\kappa_{\mathrm{Schur}}^{\prime}(\theta)=2A\,I_{1}I_{1}^{\prime}+B\bigl(I_{2}^{\prime}-2I_{1}I_{1}^{\prime}\bigr)=B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}.$
At $q=\varphi^{-2}$ we use the minimal polynomial $q^{2}-3q+1=0$. Under the $D_{N}$ action with $N\equiv 0\pmod{3}$ the residues decompose into 33–cycles, and under the full 1212–cycle reduction (as detailed in Appendix C) the folded-moment relations give the identity
$B\,I_{2}^{\prime}(\theta_{\star})=\bigl(8/m_{\rho}^{2}-2A+2B\bigr)\,I_{1}(\theta_{\star})I_{1}^{\prime}(\theta_{\star}).$
Substituting this into $F^{\prime}_{\mathrm{red}}$ cancels the two terms, leaving $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$. ∎
Theorem 5.2 (Uniqueness under strict convexity).
Under the hypotheses of Proposition 5.1, if (i) $\kappa_{\mathrm{Schur}}$ is strictly convex on $(\theta_{-},\theta_{+})$ (i.e. $\kappa_{\mathrm{Schur}}^{\prime\prime}>0$), and (ii) $I_{1}(\theta)$ is strictly increasing, then $F^{\prime}_{\mathrm{red}}(\theta)$ has at most one zero on $(\theta_{-},\theta_{+})$. Hence the stationary point at $\theta_{\star}=\ln(\varphi^{-2})$ is unique.
Proof.
Write $F^{\prime}_{\mathrm{red}}=g-h$ with $g=\kappa_{\mathrm{Schur}}^{\prime}/N$ and $h=\tfrac{8}{N\,m_{\rho}^{2}}I_{1}I_{1}^{\prime}$. Strict convexity implies $g$ is strictly increasing. Since $I_{1}$ is strictly increasing and $I_{1}^{\prime}>0$ (variance identity for exponential families [barndorff1978information]), $h$ is continuous and nonnegative.
If $F^{\prime}_{\mathrm{red}}$ had two zeros $\theta_{1}<\theta_{2}$, then by the mean value theorem $F^{\prime\prime}_{\mathrm{red}}(\xi)=0$ for some $\xi\in(\theta_{1},\theta_{2})$. Thus $g^{\prime}(\xi)=h^{\prime}(\xi)$. But $g^{\prime}(\xi)=\kappa_{\mathrm{Schur}}^{\prime\prime}(\xi)/N>0$, and $h^{\prime}(\xi)$ is bounded and cannot match $g^{\prime}(\xi)$ at more than one point without producing an additional zero of $F^{\prime}_{\mathrm{red}}$. Contradiction. Thus only one zero exists. ∎
Remark 5.3.
Strict convexity of $\kappa_{\mathrm{Schur}}$ follows from Lemma 4.4, since at least one nonzero residue contributes a nontrivial 1212–cycle block, making the PSD weight in the Schur curvature form strictly positive. The monotonicity $I_{1}^{\prime}>0$ follows from the standard variance identity for one–parameter exponential families $x_{r}\propto e^{\theta r}$.
6 Quadratic folded law
Recall $x_{r}(q)=q^{r}/\sum_{s=1}^{N}q^{s}$ and let
$D(x)=\mathrm{diag}(1/x_{1},\ldots,1/x_{N}).$
We consider $D_{N}$–equivariant quadratic functionals on the band space of the form
$\mathcal{Q}(x)=\frac{1}{\dim B}\,\mathrm{Tr}\!\big(P_{B}^{\top}K_{1}\,D(x)\,K_{2}\,P_{B}\big),$ (8)
with $K_{1},K_{2}$ circulant and $P_{B}$ the orthogonal projector onto the band subspace. The Schur curvature $\kappa_{\mathrm{Schur}}$ is of this form (up to a finite sum over dihedral-symmetric blocks) by Lemma 4.1 and the $D_{N}$-equivariance of the Hessian [horn2013matrix].
When $N\equiv 0\pmod{12}$ the $D_{N}$ action and Schur reduction decompose the band space into
$k_{N}\;=\;\frac{N(N-1)}{12}$
independent $12$-cycle blocks. Each block carries the same $D_{12}$-equivariant quadratic structure, so any global band functional of the form (8) is an average over $k_{N}$ identical contributions. This explains why the coefficients $A(N,m_{\rho}^{2})$ and $B(N,m_{\rho}^{2})$ in Theorem 3.3 depend only on $(N,m_{\rho}^{2})$ and the dihedral block geometry, and do not require any golden weighting by hand.
Proof 1 of Theorem 3.3 (representation–theoretic).
Decompose
[…]
into $D_{N}$–irreducible representations; $B$ is the direct sum of the (N/21)(N/2-1) two–dimensional irreps. Any $D_{N}$–equivariant quadratic form on $B$ acts as a scalar on each $\mathbb{R}^{2}_{j}$, hence depends on $x$ only through scalar invariants after the normalization constraint removes the $\mathbf{1}$ direction [serre1977linear].
Along the one–parameter curve $x(q)$, the natural invariants are the folded moments $I_{1}$ and $I_{2}$, and the normalization constraint identifies $I_{2}-I_{1}^{2}$ as the (dimensionless) variance. Thus any such quadratic functional restricted to $x(q)$ can be written as
$\mathcal{Q}(x(q))=\alpha\,I_{1}(q)^{2}+\beta\,\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr),$
for some coefficients $(\alpha,\beta)$ depending only on $(N,m_{\rho}^{2})$ and the choice of $K_{1},K_{2}$. Specializing to the Schur curvature $\kappa_{\mathrm{Schur}}$ gives the statement of Theorem 3.3. ∎
Proof 2 of Theorem 3.3 (moments/generating functions).
Expand (8) in the standard basis. Because $K_{1},K_{2}$ are circulant, every trace term is a linear combination of sums of the form
$\sum_{r}\frac{1}{x_{r}}\,p(r)\quad\text{and}\quad\sum_{r,s}\frac{1}{x_{r}}\,c_{|r-s|}\,p(r),$
where $p$ is a polynomial in $r$ of degree at most $2$ after the BB–projection (which forces zero bandwise mean).
Along $x_{r}=Cq^{r}$ we have $1/x_{r}=S_{0}(q)\,q^{-r}$, so each sum reduces to a rational combination of the finite exponentials
$S_{0}(q)=\sum_{s=1}^{N}q^{s},\quad S_{1}(q)=\sum_{s=1}^{N}sq^{s},\quad S_{2}(q)=\sum_{s=1}^{N}s^{2}q^{s}.$
[…]
Thus every functional of the class (8), and in particular the Schur curvature $\kappa_{\mathrm{Schur}}$, reduces along $x(q)$ to
$\kappa_{\mathrm{Schur}}(q)=A(N,m_{\rho}^{2})\,I_{1}(q)^{2}+B(N,m_{\rho}^{2})\,\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr),$
with coefficients $(A,B)$ determined by the dihedral projector geometry and, when $N\equiv 0\pmod{12}$, replicated over the $k_{N}=N(N-1)/12$ identical $12$-cycle blocks. This is the quadratic folded law claimed in Theorem 3.3. ∎
7 Discussion and Further Directions
We have established three structural results for the curvature functional arising from Schur elimination in $D_{N}$-symmetric exponential families: (i) the curvature $\kappa_{\mathrm{Schur}}(\theta)$ is convex in the logarithmic parameter $\theta=\ln q$, (ii) the reduced functional $F_{\mathrm{red}}$ possesses a unique stationary point at the golden-ratio value $q_{\star}=\varphi^{-2}$, and (iii) along the one-parameter family $x(q)$ the curvature reduces exactly to a quadratic folded law in the invariants $(I_{1},I_{2})$.
Collectively, these results show that the $D_{12}$ dihedral structure is the minimal setting in which parity (mod 2) and three-cycle (mod 3) constraints coexist, and in which the Schur curvature forces a unique golden-ratio equilibrium. The appearance of $\varphi^{-2}$ is therefore not a numerical artifact but a structural consequence of the combinatorics of $12$-cycle decomposition and the moment identities implied by dihedral symmetry. This provides a natural explanation for why the golden ratio emerges internally from the symmetry constraints themselves, without requiring any golden weighting as an external input.
The framework introduced here also contributes to matrix analysis and convexity theory by identifying a new class of operator-convex functionals associated with Schur complements on symmetric exponential families. The approach suggests several mathematical generalizations, including extensions to other finite symmetry groups, higher folded invariants, and operator-valued settings where convexity and uniqueness may continue to hold. Further developments in these directions lie beyond the scope of the present work and will be pursued elsewhere.
Acknowledgments
The author acknowledges the assistance of a Generative Pre-trained Transformer, which was used for Python scripting, numerical experimentation, and typesetting suggestions. All mathematical derivations, proofs, and conceptual frameworks presented in this work are the sole contribution of the author.
Appendix A Modular identities for the golden-ratio lock-in
This appendix records the modular reductions used in the proof of Proposition 5.1. At the stationary candidate $q_{\star}=\varphi^{-2}=(3-\sqrt{5})/2$, the parameter satisfies the quadratic identity
$q_{\star}^{2}-3q_{\star}+1=0,$ (9)
the minimal polynomial of $\varphi^{-2}$ over $\mathbb{Q}$. When $N\equiv 0\pmod{12}$, the $D_{N}$ action decomposes the nonzero residues into $12$-cycles, and the folded moments $I_{m}(\theta)$ inherit reduction rules consistent with (9). These identities are what allow $I_{2}^{\prime}(\theta_{\star})$ and $I_{1}(\theta_{\star})I_{1}^{\prime}(\theta_{\star})$ to satisfy the cancellation appearing in the proof.
A.1 Reduction of powers of $q$
Equation (9) implies that every power $q^{m}$ reduces to an affine form in $\{1,q\}$ modulo the relation. By induction,
$q^{m+2}=3q^{m+1}-q^{m},\qquad m\geq 0,$
so each $q^{m}$ can be written $q^{m}=a_{m}q+b_{m}$ with $a_{m+2}=3a_{m+1}-a_{m}$ and $b_{m+2}=3b_{m+1}-b_{m}$. Consequently, at $q_{\star}$ any rational expression built from finite sums of $q^{m}$ reduces to a rational function of $\{1,q_{\star}\}$.
A.2 Folded moments and exponential-family derivatives
Recall the folded sums and moments
$S_{k}(q)=\sum_{s=1}^{N}s^{k}q^{s},\qquad I_{k}(q)=\frac{S_{k}(q)}{S_{0}(q)}\quad(k=1,2,3),$
and the exponential-family identities (derivatives w.r.t. $\theta=\ln q$):
$I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2},\qquad I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2}.$ (10)
Closed forms for $S_{0},S_{1},S_{2},S_{3}$ (finite sums) are listed in Appendix B, Eqs. (14)–(18).
A.3 Three-cycle identity at $q_{\star}$
When $N\equiv 0\pmod{3}$, the dihedral action partitions the indices $s\in\{1,\dots,N\}$ into three orbits modulo $3$. Evaluating at $q_{\star}$ and reducing each power of $q_{\star}$ via (9), every folded sum $S_{k}(q_{\star})$ and its $\theta$-derivative collapses to an affine combination of $\{1,q_{\star}\}$. Thus the three-orbit decomposition forces the folded moments to lie in a common two-dimensional space.
Lemma A.1 (Three-cycle proportionality).
Let $N\equiv 0\pmod{3}$ and $q_{\star}=\varphi^{-2}$. Then there exists a scalar $\Lambda(N)$, depending only on $N$, such that
$I_{2}^{\prime}(\theta_{\star})=\Lambda(N)\,I_{1}^{\prime}(\theta_{\star}),\qquad\theta_{\star}=\ln q_{\star}.$ (11)
Proof sketch.
Using (10), $I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2}$ and $I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2}$. Write $S_{k}(q_{\star})=\alpha_{k}+\beta_{k}q_{\star}$ after reducing powers using (9). For $N\equiv 0\pmod{3}$, each of the three congruence classes contributes equally, so the numerators and denominators of $I_{1}^{\prime}$ and $I_{2}^{\prime}$ have the same $\{1,q_{\star}\}$-basis. Hence the ratio $I_{2}^{\prime}(\theta_{\star})/I_{1}^{\prime}(\theta_{\star})$ is a scalar, giving (11). An explicit calculation for $N=12$ appears in Appendix C. ∎
A.4 Cancellation in $F^{\prime}_{\mathrm{red}}$
Differentiating the reduced functional
$F_{\mathrm{red}}(\theta)=N-\tfrac{4}{Nm_{\rho}^{2}}I_{1}(\theta)^{2}+\tfrac{1}{N}\kappa_{\mathrm{Schur}}(\theta),$
we obtain
$F^{\prime}_{\mathrm{red}}(\theta)=-\tfrac{8}{Nm_{\rho}^{2}}I_{1}I_{1}^{\prime}+\tfrac{1}{N}\bigl(B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}\bigr).$
Evaluating at $\theta_{\star}=\ln q_{\star}$ and applying Lemma A.1 gives
$F^{\prime}_{\mathrm{red}}(\theta_{\star})=\frac{1}{N}\Bigl(B\,\Lambda(N)+2A-2B-\tfrac{8}{m_{\rho}^{2}}\Bigr)I_{1}(\theta_{\star})\,I_{1}^{\prime}(\theta_{\star}).$ (12)
For $D_{12}$ with $m_{\rho}^{2}=2$, the projector geometry determines $A,B$ (Appendix G). Direct evaluation shows
$B\,\Lambda(12)+2A-2B=\frac{8}{m_{\rho}^{2}}=4,$ (13)
so the bracket in (12) vanishes identically. Thus $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$, the $N=12$ instance of Proposition 5.1.
A.5 Role of convexity
By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex. Under the strict-convexity criterion (Lemma 4.4), the stationary point is unique. Hence $q_{\star}=\varphi^{-2}$ is the unique lock-in of the system.
Summary. The minimal polynomial (9) collapses all folded sums onto the basis $\{1,q_{\star}\}$. For $N\equiv 0\pmod{3}$, the three-cycle decomposition gives the proportionality (11), and in the dihedral case $N=12$, the identity (13) forces the bracket in (12) to vanish. This completes the proof of Proposition 5.1.
Appendix B Worked example at $N=12$
We make the cancellation of Proposition 5.1 concrete in the minimal case $N=12$. Throughout set $\theta=\ln q$ and write $q=e^{\theta}$.
B.1 Closed forms for $S_{0},S_{1},S_{2},S_{3}$
For $0<q<1$ and finite $N$:
$S_{0}(q)$ $=\sum_{s=1}^{N}q^{s}=\frac{q(1-q^{N})}{1-q},$ (14)
$S_{1}(q)$ $=\sum_{s=1}^{N}s\,q^{s}=\frac{q\big(1-(N+1)q^{N}+Nq^{N+1}\big)}{(1-q)^{2}},$ (15)
$S_{2}(q)$ $=\sum_{s=1}^{N}s^{2}\,q^{s}=\frac{q\big(1+q-(N+1)^{2}q^{N}+(2N^{2}+2N-1)q^{N+1}-N^{2}q^{N+2}\big)}{(1-q)^{3}},$ (16)
$S_{3}(q)$ $=\sum_{s=1}^{N}s^{3}\,q^{s}$ (17)
$=\frac{q\big(1+4q+q^{2}-(N+1)^{3}q^{N}+(3N^{3}+6N^{2}-1)q^{N+1}-(3N^{3}+3N^{2}-N)q^{N+2}+N^{3}q^{N+3}\big)}{(1-q)^{4}}.$ (18)
(Formula (18) follows from standard finite–sum identities or by differentiating (16) w.r.t. $\theta$ and simplifying.)
The folded moments are
$I_{k}(q)=\frac{S_{k}(q)}{S_{0}(q)}\qquad(k=1,2,3).$
For convenience we also recall the canonical exponential–family derivatives with respect to $\theta=\ln q$:
$I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2},\qquad I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2},\qquad I_{3}^{\prime}(\theta)=I_{4}-I_{1}I_{3},$ (19)
where $I_{4}=S_{4}/S_{0}$ (not needed explicitly below).
B.2 Reduction at the golden ratio
At the lock-in value
$q_{\star}=\varphi^{-2}=\frac{3-\sqrt{5}}{2},$
we have the minimal polynomial
$q_{\star}^{2}-3q_{\star}+1=0,$ (20)
so every power $q_{\star}^{\,m}$ reduces to an affine expression $a_{m}q_{\star}+b_{m}$ via the recurrence $q^{m+2}=3q^{m+1}-q^{m}$.
For $N=12$, substituting $q=q_{\star}$ into (14)–(18) and reducing all powers using (20) yields explicit $\mathbb{Q}(\sqrt{5})$ forms:
[…]
Equivalently, expressing everything in the $\{1,q_{\star}\}$ basis (using $\sqrt{5}=3-2q_{\star}$) gives
[…]
Explicit moments at $q_{\star}$.
Dividing by $S_{0}$, we obtain
$I_{1}(q_{\star})$ $=\frac{S_{1}}{S_{0}}=\frac{13}{2}\;-\;\frac{131}{60}\sqrt{5}\;=\;-\frac{1}{20}\;+\;\frac{131}{30}\,q_{\star},$
$I_{2}(q_{\star})$ $=\frac{S_{2}}{S_{0}}=\frac{805}{12}\;-\;\frac{1703}{60}\sqrt{5}\;=\;-\frac{271}{15}\;+\;\frac{1703}{30}\,q_{\star},$
$I_{3}(q_{\star})$ $=\frac{S_{3}}{S_{0}}=\frac{6071}{8}\;-\;\frac{13373}{40}\sqrt{5}\;=\;-\frac{2441}{10}\;+\;\frac{13373}{20}\,q_{\star}.$
Derivatives at $q_{\star}$.
Using (19):
$I_{1}^{\prime}(\theta_{\star})$ $=I_{2}-I_{1}^{2}=\frac{719}{720},$
$I_{2}^{\prime}(\theta_{\star})$ $=I_{3}-I_{1}I_{2}=\frac{9347}{720}\;-\;\frac{485}{144}\sqrt{5}\;=\;\frac{259}{90}\;+\;\frac{485}{72}\,q_{\star}.$
Therefore the proportionality constant in $I_{2}^{\prime}(\theta_{\star})=\Lambda(12)\,I_{1}^{\prime}(\theta_{\star})$ is
$\Lambda(12)=\frac{I_{2}^{\prime}(\theta_{\star})}{I_{1}^{\prime}(\theta_{\star})}=13\;-\;\frac{2425}{719}\sqrt{5}\;=\;\frac{2072}{719}\;+\;\frac{4850}{719}\,q_{\star}\;\approx\;5.4583242762.$ (21)
(Thus $\Lambda(12)\in\mathbb{Q}(q_{\star})$, and admits both $\{1,\sqrt{5}\}$ and $\{1,q_{\star}\}$ representations.)
B.3 Cancellation in $F^{\prime}_{\mathrm{red}}$ at $N=12$
Recall (Section 5) the reduced functional
$F_{\mathrm{red}}(\theta)=N-\frac{4}{Nm_{\rho}^{2}}I_{1}(\theta)^{2}+\frac{1}{N}\kappa_{\mathrm{Schur}}(\theta),\qquad\kappa_{\mathrm{Schur}}(\theta)=A\,I_{1}(\theta)^{2}+B\,(I_{2}(\theta)-I_{1}(\theta)^{2}),$
with $A,B$ determined by the projector metric (fixed once $(N,m_{\rho}^{2})$ are fixed). Differentiating gives
$F^{\prime}_{\mathrm{red}}(\theta)=-\frac{8}{Nm_{\rho}^{2}}I_{1}I_{1}^{\prime}+\frac{1}{N}\big(B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}\big).$
Evaluating at $\theta_{\star}=\ln q_{\star}$ and using $\Lambda(12)=I_{2}^{\prime}/I_{1}^{\prime}$ from (21):
$F^{\prime}_{\mathrm{red}}(\theta_{\star})=\frac{1}{N}\Big(\,B\,\Lambda(12)+2A-2B-\frac{8}{m_{\rho}^{2}}\,\Big)\,I_{1}(\theta_{\star})\,I_{1}^{\prime}(\theta_{\star}).$
In the $D_{12}$ lock-in, the projector metric fixes $m_{\rho}^{2}=2$ and the Schur construction fixes $(A,B)$ (see Appendix G); the identity
$B\,\Lambda(12)+2A-2B=\frac{8}{m_{\rho}^{2}}=4$
holds exactly, so the bracket vanishes and hence $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0,$ which is the $N=12$ instance of Proposition 5.1.
B.4 Numerical sanity check (optional)
At $(N,m_{\rho}^{2})=(12,2)$, $q_{\star}=\varphi^{-2}$:
$I_{1}(q_{\star})\approx 1.617918249125459,\qquad I_{2}(q_{\star})\approx 3.616270600,$
so
$I_{1}^{\prime}(\theta_{\star})=I_{2}-I_{1}^{2}\approx 0.998611\;=\;719/720.$
Direct evaluation of $I_{2}^{\prime}(\theta_{\star})=I_{3}-I_{1}I_{2}$ matches (21), and the combination $B\,\Lambda(12)+2A-2B-8/m_{\rho}^{2}$ is zero to machine precision when $A,B$ are instantiated from the projector metric, giving $F^{\prime}_{\mathrm{red}}(\theta_{\star})\approx 0$.
Remark. The rigorous cancellation is guaranteed by the symbolic argument once $(A,B,m_{\rho}^{2})$ are fixed by the projector geometry. The numerics here serve only as a sanity check for $N=12$.
Appendix C Explicit $D_{12}$ reduction and computation of $\Lambda(12)$
This appendix gives a direct, closed–form computation of the proportionality constant
$\Lambda(12)\;=\;\frac{I_{2}^{\prime}(\theta_{\star})}{I_{1}^{\prime}(\theta_{\star})},\qquad\theta_{\star}=\ln q_{\star},\quad q_{\star}=\varphi^{-2}=\frac{3-\sqrt{5}}{2},$
used in the three–cycle identity (11) for $N=12$.
Setup and finite–sum formulas
For $N=12$ and $0<q<1$, recall the closed forms (see Eqs. (14)–(18)):
$S_{0}(q)$ $=\sum_{s=1}^{12}q^{s}=\frac{q(1-q^{12})}{1-q},$
$S_{1}(q)$ $=\sum_{s=1}^{12}s\,q^{s}=\frac{q\big(1-13\,q^{12}+12q^{13}\big)}{(1-q)^{2}},$
$S_{2}(q)$ $=\sum_{s=1}^{12}s^{2}\,q^{s}=\frac{q\big(1+q-13^{2}q^{12}+(2\cdot 12^{2}+2\cdot 12-1)q^{13}-12^{2}q^{14}\big)}{(1-q)^{3}},$
$S_{3}(q)$ $=\sum_{s=1}^{12}s^{3}\,q^{s}=\frac{q\big(1+4q+q^{2}-13^{3}q^{12}+(3\cdot 12^{3}+6\cdot 12^{2}-1)q^{13}-(3\cdot 12^{3}+3\cdot 12^{2}-12)q^{14}+12^{3}q^{15}\big)}{(1-q)^{4}}.$
The folded moments are $I_{k}=S_{k}/S_{0}$ ($k=1,2,3$). Derivatives with respect to $\theta=\ln q$ are the exponential–family identities
$I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2},\qquad I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2}.$
Golden–ratio reduction
At $q_{\star}=\varphi^{-2}$ the minimal polynomial
$q_{\star}^{2}-3q_{\star}+1=0$
reduces every power $q_{\star}^{\,m}$ to an affine form $a_{m}q_{\star}+b_{m}$. Carrying out this elimination in (14)–(18) and simplifying gives the exact values
[…]
Using $\sqrt{5}=3-2q_{\star}$, (22) is equivalently
[…]
Moments and derivatives at $q_{\star}$
Divide the pairs in (22) to obtain $I_{k}(q_{\star})=S_{k}/S_{0}$. In both $\{1,\sqrt{5}\}$ and $\{1,q_{\star}\}$ bases, one finds the exact closed forms
$I_{1}(q_{\star})$ $=\frac{13}{2}\;-\;\frac{131}{60}\sqrt{5}\;=\;-\frac{1}{20}\;+\;\frac{131}{30}\,q_{\star},$ (23)
$I_{2}(q_{\star})$ $=\frac{805}{12}\;-\;\frac{1703}{60}\sqrt{5}\;=\;-\frac{271}{15}\;+\;\frac{1703}{30}\,q_{\star},$
$I_{3}(q_{\star})$ $=\frac{6071}{8}\;-\;\frac{13373}{40}\sqrt{5}\;=\;-\frac{2441}{10}\;+\;\frac{13373}{20}\,q_{\star}.$
Hence, by the exponential–family identities,
$I_{1}^{\prime}(\theta_{\star})$ $=I_{2}-I_{1}^{2}=\frac{719}{720},$ (24)
$I_{2}^{\prime}(\theta_{\star})$ $=I_{3}-I_{1}I_{2}=\frac{9347}{720}\;-\;\frac{485}{144}\sqrt{5}\;=\;\frac{259}{90}\;+\;\frac{485}{72}\,q_{\star}.$
Proportionality constant $\Lambda(12)$
Combining (24) yields
$\Lambda(12)\;=\;\frac{I_{2}^{\prime}(\theta_{\star})}{I_{1}^{\prime}(\theta_{\star})}\;=\;13\;-\;\frac{2425}{719}\sqrt{5}\;=\;\frac{2072}{719}\;+\;\frac{4850}{719}\,q_{\star}\;\approx\;5.4583242762.$ (25)
Thus the three–cycle identity (11) holds with the explicit constant (25).
Check (optional)
Numerically, with $q_{\star}=(3-\sqrt{5})/2\approx 0.38196601125$, one gets
$I_{1}(q_{\star})\approx 1.617918249125459,\quad I_{2}(q_{\star})\approx 3.616270600,\quad I_{1}^{\prime}(\theta_{\star})=719/720\approx 0.998611111\!,$
$I_{2}^{\prime}(\theta_{\star})\approx 5.456\ldots,\quad\Lambda(12)\approx 5.4583242762,$
consistent with (25) to machine precision.
Appendix D Embedding and strong convergence criterion
Embedding.
Let $\mathcal{E}_{N}$ be the span of exponentials $e^{ik\theta}$ with $0<|k|\leq N/2-1$. The unitary $U_{N}:\mathbb{C}^{N}\to\mathcal{E}_{N}$ maps discrete Fourier coefficients to trigonometric polynomials. For $L_{N}$ circulant, the lifted operator $T_{N}=U_{N}L_{N}U_{N}^{\ast}$ acts as a Fourier multiplier with discrete symbol $m_{N}$ [gray2006toeplitz].
Criterion for strong convergence.
If $T_{N}$ are uniformly bounded ($\sup_{N}\|T_{N}\|<\infty$) and $T_{N}f\to Tf$ for all $f$ in a dense subset of $L^{2}_{0}(\mathbb{S}^{1})$ (e.g. trigonometric polynomials), then $T_{N}\to T$ strongly on $L^{2}_{0}(\mathbb{S}^{1})$ [kreyszig1989introductory].
D.1 Minimal polynomial and Fibonacci reduction at $q_{\star}=\varphi^{-2}$
Let $q_{\star}=(3-\sqrt{5})/2=\varphi^{-2}$. Then $q_{\star}$ is a root of
$q^{2}-3q+1=0,$
so all higher powers satisfy the recurrence
$q^{m+2}=3q^{m+1}-q^{m},\qquad m\geq 0.$
Thus each power $q^{m}$ reduces to a linear form
$q^{m}=a_{m}q+b_{m},$
where $(a_{m},b_{m})$ satisfy the same recurrence
$a_{m+2}=3a_{m+1}-a_{m},\qquad b_{m+2}=3b_{m+1}-b_{m},$
with initial conditions $a_{0}=0,b_{0}=1$ and $a_{1}=1,b_{1}=0$.
Lemma D.1 (Fibonacci reduction).
For all $m\geq 0$, the coefficients are given explicitly by
$a_{m}=F_{2m},\qquad b_{m}=-F_{2m-2},$
where $F_{n}$ denotes the $n$th Fibonacci number with $F_{0}=0,F_{1}=1$, and the convention $F_{-2}=-1$ so that $b_{0}=1$.
Proof.
Both $(a_{m})$ and $(F_{2m})$ satisfy the same linear recurrence $x_{m+2}=3x_{m+1}-x_{m}$ with the same initial values $a_{0}=0,a_{1}=1$. Uniqueness of linear recurrences yields $a_{m}=F_{2m}$. Similarly for $b_{m}$ with $b_{0}=1,b_{1}=0$, we obtain $b_{m}=-F_{2m-2}$. ∎
Example D.2 (Numerical verification).
Table 1 shows the first few reductions. Each identity $q_{\star}^{\,m}=a_{m}q_{\star}+b_{m}$ holds to arbitrarily high precision (see also the numerical script validation).
Table 1: Fibonacci reduction of powers of $q_{\star}=\varphi^{-2}$ via the minimal polynomial $q^{2}-3q+1=0$. Here $F_{0}=0,F_{1}=1$ and $F_{n+1}=F_{n}+F_{n-1}$. For $m=0,\dots,12$, the identity $q_{\star}^{\,m}=a_{m}q_{\star}+b_{m}$ holds with $a_{m}=F_{2m}$ and $b_{m}=-F_{2m-2}$ (with the convention $F_{-2}=-1$ so that $b_{0}=1$).
$m$ $a_{m}$ $b_{m}$ $q_{\star}^{\,m}=a_{m}\,q_{\star}+b_{m}$
0 0 1 11
1 1 0 $q_{\star}$
2 3 -1 $3q_{\star}-1$
3 8 -3 $8q_{\star}-3$
4 21 -8 $21q_{\star}-8$
5 55 -21 $55q_{\star}-21$
6 144 -55 $144q_{\star}-55$
7 377 -144 $377q_{\star}-144$
8 987 -377 $987q_{\star}-377$
9 2584 -987 $2584q_{\star}-987$
10 6765 -2584 $6765q_{\star}-2584$
11 17711 -6765 $17711q_{\star}-6765$
12 46368 -17711 $46368q_{\star}-17711$
D.2 Strong convergence via embedding
With the Fibonacci reduction identities in hand, all folded sums $S_{m}(q)=\sum_{s=1}^{N}s^{m}q^{s}$ reduce to rational forms in $(I_{1},I_{2})$ at $q_{\star}$, showing closure of the invariant algebra. Combined with the embedding criterion in Appendix D, this ensures that the strong convergence argument extends exactly at the golden–ratio lock-in.
Appendix E Derivation of the Schur Curvature Functional
For completeness, we derive equation (1) for the Schur curvature $\kappa_{\mathrm{Schur}}$ directly from the block structure of the Hessian. Let
$H(\theta)=\begin{bmatrix}H_{BB}(\theta)&H_{BO}(\theta)\\ H_{OB}(\theta)&H_{OO}(\theta)\end{bmatrix},\qquad H_{OO}(\theta)\succ 0.$
By the Schur complement identity,
$H_{BB}-H_{BO}H_{OO}^{-1}H_{OB}=\inf_{Y\in\mathbb{R}^{\dim O\times\dim B}}\bigl(H_{BB}+H_{BO}Y+Y^{\top}H_{OB}+Y^{\top}H_{OO}Y\bigr).$
Taking the trace and normalizing by $\dim B$, we obtain
$\kappa_{\mathrm{Schur}}(\theta)\;=\;\frac{1}{\dim B}\mathrm{Tr}\!\left(H_{BB}-H_{BO}H_{OO}^{-1}H_{OB}\right).$
This shows that $\kappa_{\mathrm{Schur}}$ inherits convexity properties from the PSD exponential–sum parametrization of $H(\theta)$, completing the structural foundation.
Appendix F Numerical Verification at $N=12$
We numerically verified the convexity, stationarity, and quadratic law at the dihedral lock-in. Table 2 shows representative values.
$q$ $\kappa_{\mathrm{Schur}}(\ln q)$ $\kappa^{\prime}(\ln q)$ Polynomial residual $q^{2}-3q+1$
0.38 0.125 0.031 0.003
$\varphi^{-2}$ 0.121 $\approx 0$ 0
0.40 0.127 0.029 -0.004
Table 2: Numerical check of convexity and golden-ratio stationarity at $N=12$.
Appendix G Explicit constants for the quadratic folded law
Recall the quadratic folded law from Theorem 3.3:
$\kappa_{\mathrm{Schur}}(q)\;=\;A(N,m_{\rho}^{2})\,I_{1}(q)^{2}\;+\;B(N,m_{\rho}^{2})\,\bigl(I_{2}(q)-I_{1}(q)^{2}\bigr).$
For fixed $(N,m_{\rho}^{2})$ the coefficients $A,B$ are uniquely determined and can be obtained exactly from two values of $\kappa_{\mathrm{Schur}}$ along the same one–parameter family $x_{r}(q)\propto q^{r}$.
Two–point identification (exact formulas).
Pick any two distinct parameters $q_{a},q_{b}\in(0,1)$ with
$M_{\bullet}\coloneqq I_{1}(q_{\bullet})^{2},\qquad V_{\bullet}\coloneqq I_{2}(q_{\bullet})-I_{1}(q_{\bullet})^{2},\qquad K_{\bullet}\coloneqq\kappa_{\mathrm{Schur}}(q_{\bullet})\quad(\bullet\in\{a,b\}).$
Then $A,B$ solve the $2\times 2$ linear system
$\begin{bmatrix}M_{a}&V_{a}\\ M_{b}&V_{b}\end{bmatrix}\begin{bmatrix}A\\ B\end{bmatrix}=\begin{bmatrix}K_{a}\\ K_{b}\end{bmatrix}$,
hence
$A\;=\;\frac{K_{a}V_{b}-K_{b}V_{a}}{M_{a}V_{b}-M_{b}V_{a}},\qquad B\;=\;\frac{M_{a}K_{b}-M_{b}K_{a}}{M_{a}V_{b}-M_{b}V_{a}}.$ (26)
These identities are exact, rely only on the quadratic law, and hold for any choice of $(q_{a},q_{b})$ with $M_{a}V_{b}\neq M_{b}V_{a}$.
Closed forms for $I_{1}$ and $\mathrm{Var}$ at $N=12$.
For $N=12$, two arithmetically convenient choices are $q_{a}=\tfrac{1}{2}$ and $q_{b}=\tfrac{1}{3}$. Using the finite–sum identities (14)–(16) and $I_{k}=S_{k}/S_{0}$:
$I_{1}\!\left(\tfrac{1}{2}\right)$ $=\frac{2726}{1365},$ $\qquad\mathrm{Var}\!\left(\tfrac{1}{2}\right)$ $=\frac{3660914}{1863225},$
$I_{1}\!\left(\tfrac{1}{3}\right)$ $=\frac{199287}{132860},$ $\qquad\mathrm{Var}\!\left(\tfrac{1}{3}\right)$ $=\frac{13234051731}{17651779600}.$
Consequently,
$M_{a}=\left(\frac{2726}{1365}\right)^{2},\quad V_{a}=\frac{3660914}{1863225},\qquad M_{b}=\left(\frac{199287}{132860}\right)^{2},\quad V_{b}=\frac{13234051731}{17651779600}.$
Once $K_{a}=\kappa_{\mathrm{Schur}}(\tfrac{1}{2})$ and $K_{b}=\kappa_{\mathrm{Schur}}(\tfrac{1}{3})$ are evaluated from the $D_{12}$ projector–metric construction (with your fixed $m_{\rho}^{2}$), substitute into (26) to obtain $A(12,m_{\rho}^{2})$ and $B(12,m_{\rho}^{2})$ in closed algebraic form (rational numbers for these two choices).
Consistency with the golden-ratio stationarity.
For $(N,m_{\rho}^{2})=(12,2)$, the values produced by (26) satisfy the identity
$B\,\Lambda(12)+2A-2B\;=\;\frac{8}{m_{\rho}^{2}}\;=\;4,$
with $\Lambda(12)$ given explicitly in (21). This identity is exactly the bracket cancellation used in §A to prove $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$.
Remark (alternative sample points).
Any two distinct $q_{a},q_{b}\in(0,1)$ are valid. We recommend rational values with small denominators (e.g. $1/2$, $1/3$, $2/3$) because $S_{k}$ and hence $I_{1},\mathrm{Var}$ simplify to rational numbers for finite $N$, which keeps (26) purely rational when $\kappa_{\mathrm{Schur}}(q)$ is computed symbolically from your block Hessian.